from: 12Apr00 at 3:00pm
with: Matt, Chris, Sunil, Chanop, Jordan, Peter, David, Luke, Harley, Samer
Status on individual activities
Status of CVS repository
Peter has corrected the Sensor Sampling quantisation by delaying the ADC reading.
Chanop has ported his controller software to execute on VxWorks, Sun and IRIX.
Sunil is waiting for an upgrade to the Swing library before his applet will run.
Chris has been developing the global Makefiles.
Jordan is carrying out more WAM sensor sampling with revised software.
Matt has implemented range estimation using the Camera Calibration parameters.
Dave has been working on the 68332 Load and Go software.
Peter to generate both Tornado 1 and Tornado 2 kernels for Kambara.
Luke to add a battery voltage input to the ADC channels.
Harley to replace the epoxy in the bottom tube with Loctite.
Some items to do:
Purchase vent for shed
Paint new cart
Robotics Project at the ANU Robotic Systems Laboratory